OpenStudy (anonymous):

guys i give medalsss!!!!! helpppp please Tom has taken out a loan for college. He started paying off the loan with a first payment of $200. Each month he pays, he wants to pay back 1.2 times as the amount he paid the month before. Explain to Tom how to represent his first 30 payments in sigma notation. Then explain how to find the sum of his first 30 payments, using complete sentences. Explain why this series is convergent or divergent.

OpenStudy (anonymous):

@phi can you help me?

OpenStudy (anonymous):

@Garadon_Shagan can you help me?

OpenStudy (anonymous):

@ganeshie8 help

ganeshie8 (ganeshie8):

where exactly are you stuck ?

OpenStudy (anonymous):

I dont know how to make the stigma

ganeshie8 (ganeshie8):

start by writing out first few payments, then u wil get bored and see a pattern..

ganeshie8 (ganeshie8):

sigma notation is easy, il help you with sigma notation... but first find out the pattern (nth term )

ganeshie8 (ganeshie8):

first payment = $200 secon payment = ? third payment = ?

OpenStudy (anonymous):

ok for the first payment he will pay 200 right, for the second one 240, the third one 288

ganeshie8 (ganeshie8):

how u got 240 for second payment ?

OpenStudy (anonymous):

200*1.2

ganeshie8 (ganeshie8):

yes, how u got 288 for 3rd payment ?

OpenStudy (anonymous):

240*1.2

ganeshie8 (ganeshie8):

and 240 = 200*1.2, right ? so can we write the 3rd payment as below : (200*1.2)*1.2 = 200*1.2^2 ?

ganeshie8 (ganeshie8):

In general, \(n\)th payment can be written as : \[\large 200(1.2)^{n-1}\]

OpenStudy (anonymous):

so if wee were to solve folve for payment #6 we just plug in 6 in the n?

ganeshie8 (ganeshie8):

exactly!

ganeshie8 (ganeshie8):

plugin n=1 and convince yourself if you really get ur first payment or not

OpenStudy (anonymous):

okay but, you know the stigma has a lower and an upper part? what do i put there?

OpenStudy (anonymous):

yeahhh i got 200 for the first one

ganeshie8 (ganeshie8):

good :) sigma notation is just a lazy way of writing sum of things

ganeshie8 (ganeshie8):

for example : \[\large \sum \limits_{n=1}^4 n \] is same as : \[\large 1+2+3+4\]

OpenStudy (anonymous):

\[\sum_{}^{?}\]

OpenStudy (anonymous):

okay so for the upper it can be 30

OpenStudy (anonymous):

and the lower n=1

ganeshie8 (ganeshie8):

yep ! you're trying to add up all the payments, right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

you're correct about the upper/lower index

OpenStudy (anonymous):

:) so how do i know if it is divergent or convergent?

ganeshie8 (ganeshie8):

so whats ur sigma notation for the sum of first 30 payments ?

ganeshie8 (ganeshie8):

Sum of first 30 payments = \[\large \sum \limits_{n=1}^{30} 200(1.2)^{n-1} \] right ?

OpenStudy (anonymous):

yes, let me calculate the answer

OpenStudy (anonymous):

i got 39562.72

ganeshie8 (ganeshie8):

you need to use `geometric series` formula to find the sum

OpenStudy (anonymous):

yes i got that answer too

OpenStudy (anonymous):

so now i need to identify if it is convertgent or divergent

OpenStudy (anonymous):

I think it is diverrgent because it has a common ratio right ?

ganeshie8 (ganeshie8):

do u mean, it has a common ratio > 1, so it is divergent ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

then you would be correct :) if |r| > 1, then the series is divergent

ganeshie8 (ganeshie8):

here r = 1.2, which is clearly greater than 1, so the series is divergent

OpenStudy (anonymous):

okay i got it< thank you so much @ganeshie8 i really appreciate it

ganeshie8 (ganeshie8):

another way to think of it is : you're adding atleast $200 every month, so the overall sum NEVER approaches any finite number, it always kepps GROWING forever...

OpenStudy (anonymous):

oh okay :)